OpenStudy (himansh):

find x and y for e^z= 4i

hartnn (hartnn):

what have you tried? e^z = 4i take natural log (ln) on both sides

OpenStudy (himansh):

x=1/2ln4 bt i m nt able to find y

hartnn (hartnn):

z = x + iy so the part without 'i' will be 'x' and the part with 'i' will be y

hartnn (hartnn):

z = ln(4i) = ln 4 + ln i = x + i y and ln i = (pi/2) i so , x = ln 2 or 2 ln 2 y = pi/2.

OpenStudy (himansh):

e(z)== 4i i.e e^x(cosy i siny)= 4 i i have to solve the given equation by this method only

OpenStudy (himansh):

e(z)= 4i i.e e^x(cosy+ i siny)= 4 i i have to solve to the given equation by this method only

OpenStudy (himansh):

here e^x.cosy =0 ___(1) and e^x.siny=4____(2) eqn (1)^2+(2)^2 e^2x= 16 x=ln4

hartnn (hartnn):

ok then e^x cos y + i e^x sin y = 0 +4i comparing thee real and imaginary parts e^x cos y = 0 ..... e^x sin y = 4 but e^x cannot be 0, so cos y = 0 so y = ...?

OpenStudy (himansh):

but m nt able to find solution of y

OpenStudy (himansh):

thn y will be (2n+1) pi/2

hartnn (hartnn):

good work finding x! :)

OpenStudy (himansh):

x= ln 4 i knw tht bt my teacher told me y =(2npi + pi/2) U( [2n+1)pi +pi/2}

OpenStudy (himansh):

find x and y for e^z= 2 here he told me value of y be {(2n+1)pi} U 2npi m confused help me

hartnn (hartnn):

cos y = 0 gives y=(2n+1) pi/2 but sin y cannot be 0. it will make imaginary part = 0.

OpenStudy (himansh):

bt in d given question imaginary prt is 0 ri8

hartnn (hartnn):

not actually, 0+4i 0 is the real part, 4i is the imaginary part

OpenStudy (himansh):

y ops sry m mistake

OpenStudy (himansh):

here sin y = 1

OpenStudy (himansh):

hartnn thnks fr ur help :) :)